In a course, 80% students who attended the lectures. For those students who attended lectures, the probability of passing the exam is 0.9, whereas for those students who did not attend the lectures, the probability of passing the exam is 0.6. What is the probability of a randomly selected student in the class to pass the exam?

b) Given that the person does not use pass the exam, what is the probability that the student did not attend the lectures?
Answer #1

a)

P(pass the exam)=P(attended and pass)+P(not attended and pass)=0.8*0.9+0.2*0.6=0.84

b)

P( not attend given not pass the exam)=P(not attended and not pass)/P(not pass)=0.2*0.4/(1-0.84)=0.5